# GDGB: A Benchmark for Generative Dynamic Text-Attributed Graph Learning

### v1

@@ D.3 Results of TDGG on Graph Structural Metrics
D.4 Results of TDGG on Textual Quality Metrics
Full results on average textual quality scores under TDGG are shown in Table 16. Detailed results for each scoring criterion—Contextual Fidelity, Personality Depth, Dynamic Adaptability, Immersive Quality, and Content Richness—rated on a 1–5 scale, are available in Table 1724. Full descriptions of these criteria are provided in Section B.2. The experimental results demonstrate that node memory or reflection mechanisms consistently improve textual generation quality compared to their absence, highlighting the necessity of these components and the value of integrating structural and textual information in DyTAG generation.
@@ D.4 Results of TDGG on Textual Quality Metrics
D.5 Results of TDGG on Graph Embedding Metrics
Full results on the graph embedding metric under TDGG are shown in Table 25. Similar to the findings on textual quality, the use of node memory and reflection mechanisms significantly enhances generation quality, underscoring the critical role of integrating structural and textual information in DyTAG generation.
@@ D.5 Results of TDGG on Graph Embedding Metrics
D.6 Results of Node Retrieval
@@
Comprehensive results on the node retrieval task, including Hit@1 and Hit@10 metrics, are presented in Table 26. Detailed results for all LLM backbones and DGNN baselines are provided in Table 2729. The experimental findings reveal that DGNNs maintain a significant advantage in node retrieval performance. However, with reduced training data availability, the GAG-General model achieves superior results than DGNNs on the Sephora and IMDB datasets under a generative paradigm. This highlights the strong dependency of DGNNs on large-scale training data and their limited generalization capability in low-data scenarios.
@@
D.7 Results of Edge Classification
Full results on Precision, Recall, and F1 metric under the edge classification task are available in Table 30. Detailed results on all LLM backbones and DGNN baselines are available in Table 31 and 32. The experimental findings demonstrate that our GAG-General model achieves excellent performance on this discriminative task under the generative paradigm. This success stems from the strong correlation between edge labels and the semantic information encoded in both node texts and generated edge texts. For instance, in e-commerce recommendation scenarios, if a negative review is generated as the edge text, the corresponding edge label (e.g., low rating) aligns closely with this sentiment. In contrast, existing DGNNs struggle to effectively capture and utilize the semantic richness of node/edge texts in DyTAG datasets, resulting in significant performance gaps on edge classification tasks.
@@ D.7 Results of Edge Classification
D.8 Results of IDGG on Graph Structural Metrics
Full results on Degree MMD, Spectra MMD, $D_{k}$, $\alpha$, and power-law validity under IDGG using other three LLM backbones are available in Table 33, 34, and 35. Compared to the TDGG task, DyTAGs generated under IDGG exhibit slightly higher Degree/Spectra MMD values, primarily due to the introduction of new node generation, which increases the complexity of the modeling process. Nevertheless, over half of the generated DyTAG graphs still satisfy the power-law distribution criterion, demonstrating the robustness of the GAG-General framework in maintaining structural quality even under more challenging generation conditions.
@@ D.8 Results of IDGG on Graph Structural Metrics
D.9 Results of IDGG on Textual Quality Metrics
Full results on average textual quality scores under IDGG are shown in Table 36. Detailed results on each scoring criterion are available in Table 37-44. Similar to the findings under the TDGG task, the experimental results demonstrate that incorporating node memory or reflection mechanisms significantly enhances the textual quality of generated DyTAGs, highlighting the importance of integrating structural and textual information in DyTAG generation.
@@ D.9 Results of IDGG on Textual Quality Metrics
D.10 Results of IDGG on Graph Embedding Metrics
Full results on the graph embedding metric under IDGG are provided in Table 45, exhibiting the importance of incorporating structural and textual information in DyTAG generation.
@@ D.10 Results of IDGG on Graph Embedding Metrics
D.11 Results of VRDAG and DG-Gen
Comprehensive results on the graph structural and the graph embedding metrics under IDGG of GAG-General, VRDAG, and DG-Gen are shown in Table 46. Since VRDAG and DG-Gen do not support text generation, textual quality comparisons are not feasible. Furthermore, node/edge representations generated by VRDAG and DG-Gen are directly used as node/edge features when computing graph embeddings. The experimental results demonstrate that GAG-General significantly outperforms VRDAG and DG-Gen in both graph structural quality and graph embedding quality. Notably, VRDAG struggles to generate high-quality datasets across all eight GDGB datasets, while DG-Gen achieves marginally better performance due to its explicit control over the size of generated graphs. These findings highlight the limitations of existing dynamic graph generation models in the DyTAG generation task and underscore the need for future models to effectively balance structural fidelity, temporal dynamics, and textual richness.
@@ D.11 Results of VRDAG and DG-Gen
Datasets Models Ours VRDAG DG-Gen
Sephora Degree MMD 0.370 0.795 0.422
Spectra MMD 0.189 0.847 0.274
@@
Graph Embedding 0.661 0.011 0.228
Dianping Degree MMD 0.150 0.887 0.167
Spectra MMD 0.351 0.808 0.245
@@
Graph Embedding 0.769 0.024 0.517
WikiRevision Degree MMD 0.102 0.807 0.197
Spectra MMD 0.105 0.776 0.159
@@
Graph Embedding 0.699 0.089 0.121
WikiLife Degree MMD 0.078 0.464 0.283
Spectra MMD 0.206 0.236 0.288
@@
Graph Embedding 0.804 0.088 0.303
IMDB Degree MMD 0.145 0.882 0.373
Spectra MMD 0.379 0.710 0.537
@@
Graph Embedding 0.729 0.095 0.293
WeiboTech Degree MMD 0.212 0.867 0.211
Spectra MMD 0.172 0.778 0.311
@@
Graph Embedding 0.695 0.085 0.292
WeiboDaily Degree MMD 0.219 0.871 0.259
Spectra MMD 0.385 0.791 0.553
@@
Graph Embedding 0.638 0.108 0.294
Cora Degree MMD 0.073 0.877 0.212
Spectra MMD 0.181 0.760 0.365
@@
Graph Embedding 0.828 0.053 0.056
Appendix E Prompts
E.1 Prompt Templates of Bipartite Graphs

### v2

@@ F.2 Results of TDGG on Graph Structural Metrics
F.3 Results of TDGG on Textual Quality Metrics
Full results on average textual quality scores under TDGG are shown in Table˜21. Detailed results for each scoring criterion—Contextual Fidelity, Personality Depth, Dynamic Adaptability, Immersive Quality, and Content Richness—rated on a 1–5 scale, are available in Table˜2229. Full descriptions of these criteria are provided in Section˜D.3. The experimental results demonstrate that node memory or reflection mechanisms consistently improve textual generation quality compared to their absence, highlighting the necessity of these components and the value of integrating structural and textual information in DyTAG generation.
@@ F.3 Results of TDGG on Textual Quality Metrics
F.4 Results of TDGG on Graph Embedding Metrics
Full results on the graph embedding metric under TDGG are shown in Table˜30. Similar to the findings on textual quality, the use of node memory and reflection mechanisms significantly enhances generation quality, underscoring the critical role of integrating structural and textual information in DyTAG generation.
@@ F.4 Results of TDGG on Graph Embedding Metrics
F.5 Results of TDGG on Direct Usability in Downstream Tasks
To evaluate the direct usability of DyTAGs generated under the TDGG task, we assess their effectiveness as substitutes for original graphs in standard transductive downstream tasks (edge classification and link prediction), following the conventional experimental settings in dynamic graph learning (Yu et al., 2023; Zhang et al., 2024a; Peng et al., 2025). We employ four established dynamic graph neural network (DGNN) models—TGN (Rossi et al., 2020), CAWN (Wang et al., 2021a), GraphMixer (Cong et al., 2023), and DyGFormer (Yu et al., 2023)—to perform edge classification and link prediction on both the generated (Gen.) and original (Ori.) graphs across four datasets: Sephora, WikiLife, Cora, and WeiboTech.
For edge classification, the results in Table˜31 show consistently small performance gaps between the generated and original graphs. The absolute differences in Precision ($|\Delta P|$), Recall ($|\Delta R|$), and F1-score ($|\Delta F|$) are uniformly below 0.04 across all models and datasets, with most values under 0.02. This indicates that the generated graphs preserve sufficient structural and textual fidelity to support accurate classification. Similarly, in link prediction (transductive setting in dynamic graph learning), the Average Precision (AP) differences ($|\Delta AP|$) between using generated and original graphs are minimal, as summarized in Table˜32. The maximum observed $|\Delta AP|$ is 0.025, further confirming that the generated DyTAGs accurately capture the temporal evolution patterns of the original graphs.These results demonstrate that DyTAGs generated by our framework in the TDGG task are of high quality and can be directly deployed in downstream applications with negligible performance degradation.
Table 31: The result of edge classification performance on generated graphs from TDGG and original graphs.
Model Dataset Type Precision Recall F1 $|\Delta P|\downarrow$ $|\Delta R|\downarrow$ $|\Delta F|\downarrow$
TGN Sephora Gen. 0.556 $\pm$ 0.009 0.490 $\pm$ 0.012 0.523 $\pm$ 0.004 0.022 0.004 0.011
Ori. 0.534 $\pm$ 0.009 0.494 $\pm$ 0.036 0.512 $\pm$ 0.026
WikiLife Gen. 0.176 $\pm$ 0.024 0.299 $\pm$ 0.017 0.191 $\pm$ 0.033 0.033 0.017 0.030
Ori. 0.209 $\pm$ 0.016 0.282 $\pm$ 0.035 0.221 $\pm$ 0.010
Cora Gen. 0.219 $\pm$ 0.016 0.241 $\pm$ 0.004 0.206 $\pm$ 0.027 0.014 0.009 0.016
Ori. 0.233 $\pm$ 0.021 0.232 $\pm$ 0.017 0.190 $\pm$ 0.033
WeiboTech Gen. 0.781 $\pm$ 0.008 0.790 $\pm$ 0.019 0.782 $\pm$ 0.012 0.004 0.019 0.026
Ori. 0.777 $\pm$ 0.009 0.771 $\pm$ 0.024 0.756 $\pm$ 0.012
CAWN Sephora Gen. 0.526 $\pm$ 0.017 0.690 $\pm$ 0.016 0.563 $\pm$ 0.014 0.005 0.006 0.007
Ori. 0.531 $\pm$ 0.029 0.696 $\pm$ 0.003 0.556 $\pm$ 0.023
WikiLife Gen. 0.190 $\pm$ 0.015 0.292 $\pm$ 0.016 0.206 $\pm$ 0.023 0.009 0.008 0.022
Ori. 0.199 $\pm$ 0.003 0.284 $\pm$ 0.033 0.228 $\pm$ 0.026
Cora Gen. 0.212 $\pm$ 0.014 0.278 $\pm$ 0.038 0.188 $\pm$ 0.032 0.010 0.009 0.016
Ori. 0.222 $\pm$ 0.020 0.269 $\pm$ 0.031 0.204 $\pm$ 0.013
WeiboTech Gen. 0.783 $\pm$ 0.018 0.787 $\pm$ 0.017 0.784 $\pm$ 0.019 0.013 0.002 0.017
Ori. 0.770 $\pm$ 0.020 0.789 $\pm$ 0.020 0.767 $\pm$ 0.031
GraphMixer Sephora Gen. 0.576 $\pm$ 0.024 0.690 $\pm$ 0.003 0.563 $\pm$ 0.018 0.023 0.005 0.012
Ori. 0.553 $\pm$ 0.020 0.685 $\pm$ 0.017 0.575 $\pm$ 0.033
WikiLife Gen. 0.227 $\pm$ 0.015 0.315 $\pm$ 0.005 0.218 $\pm$ 0.021 0.014 0.014 0.009
Ori. 0.213 $\pm$ 0.015 0.301 $\pm$ 0.013 0.209 $\pm$ 0.008
Cora Gen. 0.222 $\pm$ 0.014 0.268 $\pm$ 0.015 0.222 $\pm$ 0.021 0.013 0.008 0.002
Ori. 0.235 $\pm$ 0.009 0.260 $\pm$ 0.033 0.224 $\pm$ 0.012
WeiboTech Gen. 0.776 $\pm$ 0.018 0.787 $\pm$ 0.036 0.775 $\pm$ 0.026 0.005 0.007 0.002
Ori. 0.771 $\pm$ 0.009 0.780 $\pm$ 0.007 0.773 $\pm$ 0.006
DyGFormer Sephora Gen. 0.476 $\pm$ 0.011 0.690 $\pm$ 0.004 0.563 $\pm$ 0.029 0.007 0.008 0.003
Ori. 0.469 $\pm$ 0.025 0.698 $\pm$ 0.007 0.560 $\pm$ 0.017
WikiLife Gen. 0.242 $\pm$ 0.024 0.311 $\pm$ 0.014 0.194 $\pm$ 0.029 0.003 0.000 0.002
Ori. 0.239 $\pm$ 0.029 0.311 $\pm$ 0.020 0.196 $\pm$ 0.017
Cora Gen. 0.209 $\pm$ 0.018 0.274 $\pm$ 0.039 0.206 $\pm$ 0.023 0.010 0.024 0.007
Ori. 0.219 $\pm$ 0.021 0.250 $\pm$ 0.024 0.213 $\pm$ 0.030
WeiboTech Gen. 0.806 $\pm$ 0.010 0.803 $\pm$ 0.024 0.780 $\pm$ 0.014 0.020 0.013 0.013
Ori. 0.786 $\pm$ 0.007 0.790 $\pm$ 0.006 0.767 $\pm$ 0.006
Table 32: The results of link prediction performance (transductive) on generated graphs from TDGG and original graphs.
Model Dataset Type Average Precision $|\Delta AP|\downarrow$
TGN Sephora Gen. 0.663 $\pm$ 0.025 0.010
Ori. 0.673 $\pm$ 0.016
WikiLife Gen. 0.807 $\pm$ 0.017 0.017
Ori. 0.790 $\pm$ 0.023
Cora Gen. 0.621 $\pm$ 0.022 0.025
Ori. 0.596 $\pm$ 0.012
WeiboTech Gen. 0.859 $\pm$ 0.009 0.021
Ori. 0.838 $\pm$ 0.027
CAWN Sephora Gen. 0.731 $\pm$ 0.017 0.001
Ori. 0.730 $\pm$ 0.015
WikiLife Gen. 0.831 $\pm$ 0.003 0.019
Ori. 0.850 $\pm$ 0.023
Cora Gen. 0.616 $\pm$ 0.011 0.024
Ori. 0.592 $\pm$ 0.017
WeiboTech Gen. 0.862 $\pm$ 0.020 0.013
Ori. 0.849 $\pm$ 0.030
GraphMixer Sephora Gen. 0.729 $\pm$ 0.019 0.018
Ori. 0.711 $\pm$ 0.010
WikiLife Gen. 0.813 $\pm$ 0.002 0.018
Ori. 0.795 $\pm$ 0.030
Cora Gen. 0.626 $\pm$ 0.030 0.002
Ori. 0.624 $\pm$ 0.021
WeiboTech Gen. 0.868 $\pm$ 0.028 0.010
Ori. 0.858 $\pm$ 0.011
DyGFormer Sephora Gen. 0.631 $\pm$ 0.007 0.013
Ori. 0.644 $\pm$ 0.027
WikiLife Gen. 0.800 $\pm$ 0.022 0.004
Ori. 0.796 $\pm$ 0.012
Cora Gen. 0.824 $\pm$ 0.020 0.004
Ori. 0.828 $\pm$ 0.027
WeiboTech Gen. 0.800 $\pm$ 0.020 0.001
Ori. 0.799 $\pm$ 0.009
F.6 Results of Node Retrieval
@@
Comprehensive results on the node retrieval task, including Hit@1 and Hit@10 metrics, are presented in Table˜33. Detailed results for all LLM backbones and DGNN baselines are provided in Table˜3436. The experimental findings reveal that DGNNs maintain a significant advantage in node retrieval performance. However, with reduced training data availability, the GAG-General model achieves superior results than DGNNs on the Sephora and IMDB datasets under a generative paradigm. This highlights the strong dependency of DGNNs on large-scale training data and their limited generalization capability in low-data scenarios.
@@
F.7 Results of Edge Classification
Full results on Precision, Recall, and F1 metric under the edge classification task are available in Table˜37. Detailed results on all LLM backbones and DGNN baselines are available in Table˜38 and 39. The experimental findings demonstrate that our GAG-General model achieves excellent performance on this discriminative task under the generative paradigm. This success stems from the strong correlation between edge labels and the semantic information encoded in both node texts and generated edge texts. For instance, in e-commerce recommendation scenarios, if a negative review is generated as the edge text, the corresponding edge label (e.g., low rating) aligns closely with this sentiment. In contrast, existing DGNNs struggle to effectively capture and utilize the semantic richness of node/edge texts in DyTAG datasets, resulting in significant performance gaps on edge classification tasks.
@@ F.7 Results of Edge Classification
F.8 Results of IDGG on Graph Structural Metrics
Full results on Degree MMD, Spectra MMD, $D_{k}$, $\alpha$, and power-law validity under IDGG using other three LLM backbones are available in Table˜40, 41, and 42. Compared to the TDGG task, DyTAGs generated under IDGG exhibit slightly higher Degree/Spectra MMD values, primarily due to the introduction of new node generation, which increases the complexity of the modeling process. Nevertheless, over half of the generated DyTAG graphs still satisfy the power-law distribution criterion, demonstrating the robustness of the GAG-General framework in maintaining structural quality even under more challenging generation conditions.
@@ F.8 Results of IDGG on Graph Structural Metrics
F.9 Results of IDGG on Textual Quality Metrics
Full results on average textual quality scores under IDGG are shown in Table˜43. Detailed results on each scoring criterion are available in Table˜44-51. Similar to the findings under the TDGG task, the experimental results demonstrate that incorporating node memory or reflection mechanisms significantly enhances the textual quality of generated DyTAGs, highlighting the importance of integrating structural and textual information in DyTAG generation.
@@ F.9 Results of IDGG on Textual Quality Metrics
F.10 Results of IDGG on Graph Embedding Metrics
Full results on the graph embedding metric under IDGG are provided in Table˜52, exhibiting the importance of incorporating structural and textual information in DyTAG generation.
@@ F.10 Results of IDGG on Graph Embedding Metrics
F.11 Results of VRDAG, DG-Gen, and TIGGER-I
Comprehensive results on the graph structural and the graph embedding metrics under IDGG of GAG-General, VRDAG, DG-Gen, and TIGGER-I are shown in Table˜53. Since VRDAG (Li et al., 2024), DG-Gen (Hosseini et al., 2024), and TIGGER-I (Gupta et al., 2022) do not support text generation, textual quality comparisons are not feasible. Furthermore, node/edge representations generated by VRDAG and DG-Gen are directly used as node/edge features when computing graph embeddings. The experimental results demonstrate that GAG-General significantly outperforms VRDAG, DG-Gen, and TIGGER-I in both graph structural quality and graph embedding quality. Notably, VRDAG and TIGGER-I struggles to generate high-quality datasets across all eight GDGB datasets, while DG-Gen achieves marginally better performance due to its explicit control over the size of generated graphs. These findings highlight the limitations of existing dynamic graph generation models in the DyTAG generation task and underscore the need for future models to effectively balance structural fidelity, temporal dynamics, and textual richness.
@@ F.11 Results of VRDAG, DG-Gen, and TIGGER-I
Datasets Models Ours VRDAG DG-Gen TIGGER-I
Sephora Degree MMD $\downarrow$ 0.370 0.795 0.422 0.622
Spectra MMD $\downarrow$ 0.189 0.847 0.274 0.687
@@
Graph Embedding $\uparrow$ 0.661 0.011 0.228 0.085
Dianping Degree MMD $\downarrow$ 0.150 0.887 0.167 0.446
Spectra MMD $\downarrow$ 0.351 0.808 0.245 0.341
@@
Graph Embedding $\uparrow$ 0.769 0.024 0.517 0.580
WikiRevision Degree MMD $\downarrow$ 0.102 0.807 0.197 0.658
Spectra MMD $\downarrow$ 0.105 0.776 0.159 0.606
@@
Graph Embedding $\uparrow$ 0.699 0.089 0.121 0.105
WikiLife Degree MMD $\downarrow$ 0.078 0.464 0.283 0.418
Spectra MMD $\downarrow$ 0.206 0.236 0.288 0.254
@@
Graph Embedding $\uparrow$ 0.804 0.088 0.303 0.156
IMDB Degree MMD $\downarrow$ 0.145 0.882 0.373 0.679
Spectra MMD $\downarrow$ 0.379 0.710 0.537 0.653
@@
Graph Embedding $\uparrow$ 0.729 0.095 0.293 0.159
WeiboTech Degree MMD $\downarrow$ 0.212 0.867 0.211 0.313
Spectra MMD $\downarrow$ 0.172 0.778 0.311 0.384
@@
Graph Embedding $\uparrow$ 0.695 0.085 0.292 0.289
WeiboDaily Degree MMD $\downarrow$ 0.219 0.871 0.259 0.465
Spectra MMD $\downarrow$ 0.385 0.791 0.553 0.589
@@
Graph Embedding $\uparrow$ 0.638 0.108 0.294 0.241
Cora Degree MMD $\downarrow$ 0.073 0.877 0.212 0.372
Spectra MMD $\downarrow$ 0.181 0.760 0.365 0.389
@@
Graph Embedding $\uparrow$ 0.828 0.053 0.056 0.197
F.12 Results of IDGG on Utility in Data Augmentation for Inductive Learning
We further investigate the utility of DyTAGs generated under the IDGG task for data augmentation in inductive learning scenarios, which simulate real-world cold-start problems such as recommending new items to users. We augment the training data of the original graph with the generated DyTAG (Aug.) and evaluate performance on an inductive link prediction task, where the goal is to predict future links involving nodes not observed during training. The DGNN models and the experimental settings that we used for evaluation are identical to those in Section˜F.5.
As shown in Table˜54, augmenting training data with IDGG-generated graphs consistently improves model performance across all datasets and models. The improvement in Average Precision ($\Delta AP$) ranges from 0.027 to 0.108, with particularly notable gains on WikiLife and Cora. For example, CAWN achieves a $\Delta AP$ of 0.108 on WikiLife, and DyGFormer shows a 0.046 improvement on Cora. These results demonstrate that IDGG-generated DyTAGs provide valuable structural and textual context for unseen nodes, effectively enriching the training signal and enhancing model generalization in cold-start settings. This validates the practical value of our generative framework in improving downstream model robustness through synthetic data augmentation.
Table 54: The results of link prediction (inductive) performance with data augmentation from the graphs generated from IDGG.
Model Dataset Type Average Precision $\Delta AP\uparrow$
TGN Sephora Aug. 0.673 $\pm$ 0.010 0.037
Ori. 0.636 $\pm$ 0.002
WikiLife Aug. 0.667 $\pm$ 0.020 0.099
Ori. 0.568 $\pm$ 0.024
Cora Aug. 0.552 $\pm$ 0.030 0.034
Ori. 0.518 $\pm$ 0.019
WeiboTech Aug. 0.741 $\pm$ 0.022 0.041
Ori. 0.700 $\pm$ 0.027
CAWN Sephora Aug. 0.750 $\pm$ 0.014 0.050
Ori. 0.700 $\pm$ 0.013
WikiLife Aug. 0.741 $\pm$ 0.010 0.108
Ori. 0.633 $\pm$ 0.006
Cora Aug. 0.651 $\pm$ 0.017 0.043
Ori. 0.608 $\pm$ 0.010
WeiboTech Aug. 0.740 $\pm$ 0.009 0.033
Ori. 0.707 $\pm$ 0.009
GraphMixer Sephora Aug. 0.710 $\pm$ 0.006 0.027
Ori. 0.683 $\pm$ 0.025
WikiLife Aug. 0.696 $\pm$ 0.004 0.031
Ori. 0.665 $\pm$ 0.030
Cora Aug. 0.575 $\pm$ 0.003 0.040
Ori. 0.535 $\pm$ 0.018
WeiboTech Aug. 0.720 $\pm$ 0.006 0.058
Ori. 0.662 $\pm$ 0.029
DyGFormer Sephora Aug. 0.691 $\pm$ 0.009 0.051
Ori. 0.640 $\pm$ 0.003
WikiLife Aug. 0.734 $\pm$ 0.022 0.039
Ori. 0.695 $\pm$ 0.023
Cora Aug. 0.599 $\pm$ 0.019 0.046
Ori. 0.553 $\pm$ 0.020
WeiboTech Aug. 0.736 $\pm$ 0.008 0.040
Ori. 0.696 $\pm$ 0.016
F.13 Visualizations of the Hub Node Structures
The visualization of hub node structures in both the ground-truth and generated graphs from IDGG is presented in Figures˜10 and 11. The top-3 hub product nodes are highlighted in dark red. As described in Section˜5.3, both graphs exhibit similar hub node patterns, yet the generated graph contains hub nodes formed through inductive generation, featuring entirely distinct textual profiles.
This divergence arises because IDGG emulates real-world graph evolution dynamics, preserving structural fidelity while generating new nodes with plausible attributes that align with the underlying generative mechanisms of real systems. For example, in recommendation systems, hub nodes in the generated DyTAG may represent emerging products with high virality potential, whereas those in the ground-truth graph correspond to established bestsellers. This capability establishes DyTAG generation as a strategic tool for proactive decision-making in e-commerce and digital marketing. By identifying potential future hubs, platforms can prioritize resource allocation for product promotion, optimize advertising strategies, and anticipate market trends before they emerge in real-world data.
Figure 10: The visualization of the hub node structures in the ground-truth graph.
Figure 11: The visualization of the hub node structures in the generated graph.
F.14 Semantic-drift Analysis for IDGG Evaluation
To assess the long-term semantic consistency of generated DyTAGs in the IDGG task, we introduce a cross-snapshot semantic drift detection evaluation. Specifically, we partition the generated 10K-edge DyTAG into 10 temporal snapshots $\{\mathcal{G}_{t}\}_{t=1}^{10}$, each containing 1K newly added edges and corresponding newly generated nodes. For each snapshot $\mathcal{G}_{t}$, we compute a global semantic representation $\mathbf{h}_{t}$ by encoding all textual node and edge attributes using BERT-base-uncased(Devlin et al., 2019) and aggregating via mean pooling:
$\mathbf{h}_{t}=\frac{1}{|\mathcal{V}_{t}|+|\mathcal{E}_{t}|}\left(\sum_{v\in\mathcal{V}_{t}}\text{BERT}(v.\text{text})+\sum_{e\in\mathcal{E}_{t}}\text{BERT}(e.\text{text})\right),$ (1)
where $\mathcal{V}_{t}$ and $\mathcal{E}_{t}$ denote the sets of nodes and edges introduced in the current snapshot $t$. We then measure inter-temporal semantic stability via cosine similarity between consecutive representations:
$s_{t}=\cos(\mathbf{h}_{t},\mathbf{h}_{t+1}),\quad t=1,\dots,9.$ (2)
Thus, higher $s_{t}$ indicates greater semantic coherence over time.
On the Sephora dataset, as shown in Figure˜12, BERT-based similarity metrics reveal a gradual decline after 5K edges under GAG-General—suggesting emerging semantic drift (e.g., generation of out-of-domain items like “makeup bag”). Importantly, we further analyze the semantic drift under guided generation. As shown in Figure˜12, we find that the drift is significantly reduced when domain-specific constraints (e.g., “only generate reviews of makeup products (edge generation)/core makeup products (node generation), which are possible to be shown on Sephora platform”) are incorporated into the prompts for edge generation (Table˜59) or node generation (Table˜60), demonstrating the effectiveness of guided generation in future work.
Figure 12: Cross-snapshot semantic consistency over three LLM backbones in IDGG on Sephora.
F.15 Human-amenity Check for IDGG Evaluation
To evaluate the practical usability and naturalness of generated content, we conduct a small-scale human evaluation on the IDGG-generated DyTAGs. Five skilled annotators independently assessed 50 randomly sampled newly generated nodes and their interactions (e.g., reviews, revisions, citations, etc.). Each item was rated on a 5-point Likert scale across three dimensions: naturalness, plausibility, and consistency with platform context.
The results are shown in Table˜55. For instance, the results on Sephora with GPT as the LLM backbone show that generated nodes (users/products) achieve an average score of 3.9, while generated edges (e.g., user reviews) receive a higher average of 4.3. These scores indicate that the majority of generated content is perceived as realistic, contextually appropriate, and aligned with real-world user behavior on the target platform—supporting the human amenity of our framework in practical deployment scenarios.
Table 55: Human evaluation average scores (5-point Likert scale) for generated nodes and edges under IDGG across different LLM backbones and datasets. Higher scores indicate better naturalness, plausibility, and consistency.
DeepSeek Llama Qwen GPT
Node Edge Node Edge Node Edge Node Edge
Sephora 3.7 4.2 3.6 4.1 3.8 4.3 3.9 4.3
Dianping 3.5 4.0 4.0 4.5 3.6 4.1 3.8 4.2
WikiRevision 3.3 3.7 3.2 3.6 3.6 4.0 3.5 3.9
WikiLife 3.6 4.1 3.5 4.0 3.7 4.2 3.8 4.3
IMDB 3.8 4.3 3.3 3.8 3.5 4.0 3.6 4.1
WeiboTech 3.2 3.6 3.1 3.5 3.3 3.7 3.4 3.8
WeiboDaily 3.3 3.8 3.8 4.3 3.4 3.9 3.5 4.0
Cora 3.5 4.0 3.9 4.5 3.6 4.1 3.7 4.2
Appendix G Prompts
G.1 Prompt Templates of Bipartite Graphs